### v2

@@ Abstract
Artificial intelligence (AI) has seen remarkable advancements across various domains, including natural language processing, computer vision, autonomous vehicles, and biology. However, the rapid expansion of AI technologies has escalated the demand for more powerful computing resources. As digital computing approaches fundamental limits, neuromorphic photonics emerges as a promising platform to complement existing digital systems. In neuromorphic photonic computing, photonic devices are controlled using analog signals. This necessitates the use of digital-to-analog converters (DAC) and analog-to-digital converters (ADC) for interfacing with these devices during inference and training. However, data movement between memory and these converters in conventional von Neumann computing architectures consumes energy. To address this, analog memory co-located with photonic computing devices is proposed. This approach aims to reduce the reliance on DACs and minimize data movement to enhance compute efficiency. This paper demonstrates a monolithically integrated neuromorphic photonic circuit with co-located capacitive analog memory and compares various analog memory technologies for neuromorphic photonic computing using the MNIST dataset as a benchmark.
@@ 1.1 Motivation
Figure 1: Motivation for electro-optic processors with dynamic electro-optic analog memory (DEOAM). a) Dynamic electro-optic analog memory (DEOAM) consists of a capacitor connected to a PN junction microring resonator (MRR). The capacitor holds data on the MRR, thereby enacting a weight in the optical domain. Using DEOAM means DACs can be reused among rows of analog memory devices since the analog memory holds the signal on the MRR, which relaxes limitations on energy efficiency and bandwidth imposed by DACs. b) Conventional electro-optic processors require a dedicated SRAM-DAC for each PN junction MRR. The DAC needs to be constantly active to hold a signal on the MRR; therefore, implementing SRAM-DAC in a neuromorphic photonic processor requires a DAC for every MRR. c) If n represents the number of MRRs in the row and column of a neuromorphic photonic processor, scaling the conventional SRAM-DAC implementation incurs an n2 DAC penalty, whereas DEOAM scales by n. Write power and sampling rate for 32x32, 8-bit memory cell architectures are compared using each memory device and DACs from literature [12, 13, 14, 15, 16, 17, 18, 19, 20]. As DEOAM scales, DEOAM saves more power compared to SRAM-DAC. Details on the power calculations are described in Appendix D.
@@
Here, we propose a neuromorphic photonic processor with dynamic electro-optic analog memory (DEOAM, Figure 1a). This processor leverages the monolithic integration of nanophotonic and CMOS devices on the same silicon substrate. The proposed architecture (Figure 1a) merges the parallel information processing capabilities of photonics using wavelength-division multiplexing (WDM) with the low control complexity of crossbar arrays offered by electronics. As a proof-of-concept, we demonstrate neurosynaptic photonic weights, realized through a WDM array of tunable PN junction microring resonators (MRR) within a photonic weightbank and integrated with an array of capacitive DEOAM on the 90 nm GF9WG monolithic process. We also evaluate and compare various analog memory technologies to enhance system performance. The MNIST dataset serves as a benchmark for training and inference in our comparative analysis of these analog memory technologies, along with a discussion on tradeoffs in neural network design [24]. Our approach to integrating analog memory in neuromorphic photonic processors aims to create a flexible processor with high bandwidth, efficient training capabilities, and the ability to adapt through on-chip, online learning.
1.2 Dynamic Electro-Optic Analog Memory
@@
In a neuromorphic photonic processor with n rows and n columns of MRRs acting as weights, DEOAM reduces the number of DACs, thereby reducing power consumption. Since the conventional SRAM-DAC implementation requires each MRR to be paired with a DAC, scaling the implementation results in the number of DACs increasing by n2 (Figure 1b). In contrast, using DEOAM reduces the number of DACs below n2 (Figure 1a) because data is held directly on the photonic device via analog memory, eliminating the constant processing typically needed by DACs. This concept is similar to the functionality in modern liquid crystal active-matrix displays [25]. DACs are still required in the DEOAM architecture to load initial weights from non-volatile, digital memory before beginning inference or training and to refresh analog memory during inference. However, this approach not only simplifies control complexities by reducing the number of electrical lines needed for DAC operations but also, with longer retention times, diminishes the need for DACs to continuously drive the photonic devices, leading to further power savings as the processor scales (Figure 1c, inset). Hence, integrating DEOAM and reducing the reliance on DACs paves the way for more efficient computing.
@@ 2.1 Electro-Optic Characteristics
2.2 Device-Level Comparison of Analog Memory Technologies
[…]
2.3 System-Level Comparison of Analog Memory Technologies
To evaluate the performance of different analog memory technologies within a neural network system, a weight bank architecture, depicted in Figure 4, is emulated to assess analog memory performance during inference and training using the MNIST dataset [24].
Figure 4: The neural network emulation architecture and initial training. a) The feedforward neural network is trained using the MNIST dataset with 50000, 1-bit, 28x28 pixel images using a batch size of 64, then validated and tested using 10000 images [24]. The neural network is three layers with the input layer supporting 784 values for the 28x28 pixel images, the hidden layer supporting 50 neurons using a ReLU activation function, and the output layer supporting 10 neurons using a logarithmic softmax activation function to classify as a number. The neural network can map to the proposed neuromorphic photonic hardware that consists of 10 photonic cores with 50 rows of weight banks and 80 MRRs in each weight bank. Each core uses 80 wavelengths and 50 semiconductor optical amplifiers (SOA) to compensate for the splitting loss. b) The neural network achieves more than 95% inference accuracy after one epoch for all numbers. The specifics of the neural network architecture, performance, and modeling are detailed in Appendix E.
In Figure 5, the weight bank array is characterized for inference accuracy by comparing inference accuracy using weights trained on a GPU with full floating point accuracy and weights trained on emulated hardware with non-idealities in the analog memory and hardware. These non-idealities include analog memory control bit precision, noise, latency, and retention time, which can be generalized to any analog memory technology. These non-idealities are examined individually to assess their impact on inference accuracy. Figure 5a reveals that to achieve inference accuracies above 95%, more than 4 control bits are required for analog memory when weights are trimmed only in inference. The minimum control bit requirement increases to over 8 control bits when weights are adjusted in training and inference because lower control bit precision during training results in less accurate gradient calculations and weights, leading to reduced inference accuracy. For DEOAM, which achieves about 5 bits of precision, using 4 control bits results in an optical power change of about 383 nW and a delay of about 4 $\mu$s before the weight is registered as a different value.
@@ 2.3 System-Level Comparison of Analog Memory Technologies
Once weights become unusable due to leakage, dead-time is consumed to refresh the weight, which can be approximated by the analog memory’s write time. When examining the ratio of retention time to write time for analog memories from Table 2, this ratio (10,000 for DEOAM) is a couple orders of magnitude greater than the retention time to network latency ratio (100) required to have high inference accuracy. This means the dead-time for refreshing can be much shorter than network latency. Therefore, refreshes can occur after each image classification, thereby maintaining accurate weights for high inference accuracy without adding significant latency but at the cost of refresh power.
Figure 5: Neural network emulation results with and without non-idealities from the analog memory and system, such as control bit precision, noise, leakage, and latency. a) Minimum analog memory control bit precision to achieve more than 95% inference accuracy is 4 bits for weights trimmed only in inference and 8 bits for weights trimmed in inference and training. Inference accuracy degrades with increasing laser, SOA, PD, and TIA noise, but inference accuracy degrades b) earlier when weights are trained without noise and c) later when weights are trained with noise. d) The sweep of network latency and retention time characterizes inference accuracy for weights trained with and without leaky analog memory and network latency. Network latency and analog memory retention time constant at the point when inference accuracy degrades by 10% is plotted. When weights are trained with the effects of network latency and memory retention time, the network compensates for these effects and allows for a lower analog memory retention time constant requirement. Batch sizes of 64, 32, and 16 reveal that large batch sizes require longer retention times since weights are updated less frequently. Weights are updated more frequently in smaller batch sizes. Appendix E expands on results in d).
@@ 3 Discussion
However, scaling the neuromorphic photonic processor with analog memory presents challenges in analog memory and MRR control. Depending on the analog memory requirements, specialized circuitry may be necessary for writing to or refreshing the memory. As the system scales, strategies like sharing drive lines (as illustrated in Figure 1a) where DACs are shared along columns, can be essential in reducing power consumption. Accurate resonance control in MRRs is vital for maintaining precise optical weight representation. As the processor scales, implementing thermal control and multi-MRR control algorithms becomes increasingly critical for system stability [59, 60, 61, 62, 39, 63].
@@ 4 Conclusion
We demonstrated a proof-of-concept neuromorphic photonic circuit with DEOAM and compared its capabilities with other analog memory technologies. DEOAM reduces DACs in a neuromorphic photonic processor from n2 to n, thereby alleviating bottlenecks in bandwidth and energy consumption. By integrating analog memory on chip, neuromorphic photonic processors have the potential to achieve fast and efficient training. Since the training process occurs on chip, the neural network can readily adapt to environmental fluctuations and noise by incorporating them during training. Moreover, analog memory with on-chip backpropagation hardware can enable online training, allowing weights to be updated after each new input data. This opens up possibilities for various applications requiring recurrent events, event-based decision-making, or adaptability, such as time series forecasting. As a result, the next generation of AI processors, equipped with integrated analog memory, can help sustain the continued growth of AI.
@@ Appendix A Hardware Setup
The write time measurement setup is shown in Figure 8b and is similar to the retention time and leakage measurement setup in Figure 8a and 8c, respectively, except that the AWG is driving the analog memory and PN junction MRR while the SMU is holding the EN signal active. The write time measurement sequence is shown in Figure 3c and 3e. The AWG sends a pulse train with 5 ns edge rates to drive the analog memory and PN junction MRR, which modulates the optical response on the THRU and DROP port. The optical response is detected using the THRU port by the PD255 and measured by the DSO81304A oscilloscope. The write time measurement sequence shown in Figure 3d differs from Figure 3c and 3e because the rather than driving the analog memory and PN junction with the 5 ns edge rate pulse train, the transmission gate is driven with a pulse train that connects and disconnects the voltage source from the analog memory and PN junction. The rise time, fall time, and one time constant values are extracted from the oscilloscope time domain response.
@@ Appendix B Optical Characteristics
From the static optical characteristics of the MRRs in Figure 9a and 9b, the minimum wavelength spacing between different rings is 0.34 nm, the free spectral range (FSR) is 4.67 nm, the extinction ratio is between 15.6 dB and 19.6 dB, and the quality factor is about 50,000. From Figure 9a, optical bus power is estimated to be -6 dBm. In Figure 9c and Figure 9d, to observe the tuning efficiency of the PN junction MRRs, one MRR is biased at various voltages from 0 to 3 V (3 V is the limit for the thick oxide transistors) while the other MRRs are biased at zero voltage. The resonance wavelength and optical power of each MRR are observed at each voltage bias point to track tuning. Figure 9c reveals that as peak 3 is biased, peaks 1, 2, and 4 track similar trajectories, meaning optical crosstalk between adjacent resonances is negligible. In Figure 9d, the resonance power shift is less apparent than the resonance wavelength shift because the power shift occurs near the resonance of the ring, where the ring response’s slope is shallow. Looking at the tuning efficiency of the MRR in Figure 9e and 9f, the wavelength and power shifts are 6.22 pm/V and 0.23 dB/V, respectively.
@@ Appendix C Electro-Optic Characteristics
Appendix D Comparing Static and Dynamic Electro-Optic Analog Memory
In electronics, static random access memory (SRAM) typically refers to memory implemented as bit cells with transistors, and dynamic random access memory (DRAM) refers to memory implemented as capacitors. Here, we detail the write power calculations for static and dynamic electro-optic analog memory, SRAM-DAC and DEOAM respectively, in a neuromorphic photonic processor with n columns of MRRs and n rows of photonic weight banks, totaling to n2 MRRs and n2 memory cells. SRAM-DAC is defined as consisting of SRAM and DACs, and DEOAM is defined as DACs connected to capacitors. The power consumed by SRAM-DAC is calculated as:
$P_{SRAM-DAC}=n^{2}P_{DAC}+P_{SRAM_{dynamic}}n+P_{SRAM_{static}}n^{2}$ (1)
where PSRAM-DAC is the total SRAM-DAC array power consumption in W, and PSRAM and PDAC are the SRAM and DAC power consumption in W, respectively, which includes static and dynamic power consumption and accounts for varying bit precisions of the DAC. All surveyed DACs are demonstrated, and the SRAMs associated with each DAC are surveyed to be in the same process node [12, 13, 14, 15, 16, 17, 18, 19, 20]. The power consumed by DEOAM is calculated as:
$P_{DEOAM}=(P_{DAC}+\alpha CfV^{2})n$ (2)
where PDEOAM is the total DEOAM array power consumption, $\alpha$ is the activity factor (assumed to be 1), C is the load capacitance in F, f is the switching frequency in Hz and is half the sampling rate, and V is the switching voltage.
Appendix E Emulation Modeling
E.1 Neural Network Architecture and Performance
@@ E.2 Noise
E.3 Weight Decay and Network Latency
@@
E.4 Effects of Retention Time Constant, Network Latency, and Batch Size on Inference Accuracy

### v3

@@ Abstract
Artificial intelligence (AI) has seen remarkable advancements across various domains, including natural language processing, computer vision, autonomous vehicles, and biology. However, the rapid expansion of AI technologies has escalated the demand for more powerful computing resources. As digital computing approaches fundamental limits, neuromorphic photonics emerges as a promising platform to complement existing digital systems. In neuromorphic photonic computing, photonic devices are controlled using analog signals. This necessitates the use of digital-to-analog converters (DAC) and analog-to-digital converters (ADC) for interfacing with these devices during inference and training. However, data movement between memory and these converters in conventional von Neumann computing architectures consumes energy. To address this, analog memory co-located with photonic computing devices is proposed. This approach aims to reduce the reliance on DACs and minimize data movement to enhance compute efficiency. This paper demonstrates a monolithically integrated neuromorphic photonic circuit with co-located capacitive analog memory and analyzes analog memory specifications for neuromorphic photonic computing using the MNIST dataset as a benchmark.
@@ 1.1 Motivation
Figure 1: Motivation for electro-optic processors with dynamic electro-optic analog memory (DEOAM). a) Dynamic electro-optic analog memory (DEOAM) consists of a capacitor connected to a PN junction microring resonator (MRR). The capacitor holds data on the MRR, thereby enacting a weight in the optical domain. Using DEOAM means DACs can be reused among rows of analog memory devices since the analog memory holds the signal on the MRR, which relaxes limitations on energy efficiency and bandwidth imposed by DACs. b) Conventional electro-optic processors require a dedicated SRAM-DAC for each PN junction MRR. The DAC needs to be constantly active to hold a signal on the MRR; therefore, implementing SRAM-DAC in a neuromorphic photonic processor requires a DAC for every MRR. c) If n represents the number of MRRs in the row and column of a neuromorphic photonic processor, scaling the conventional SRAM-DAC implementation incurs an n2 DAC penalty, whereas DEOAM scales by n. Write power and sampling rate for 32x32, 8-bit memory cell architectures are compared using each memory device and DACs from literature [12, 13, 14, 15, 16, 17, 18, 19, 20]. As DEOAM scales, DEOAM saves more power compared to SRAM-DAC.
@@
Here, we propose a neuromorphic photonic processor with dynamic electro-optic analog memory (DEOAM, Figure 1a). This processor leverages the monolithic integration of nanophotonic and CMOS devices on the same silicon substrate. The proposed architecture (Figure 1a) merges the parallel information processing capabilities of photonics using wavelength-division multiplexing (WDM) with the low control complexity of crossbar arrays offered by electronics. As a proof-of-concept, we demonstrate neurosynaptic photonic weights, realized through a WDM array of tunable PN junction microring resonators (MRR) within a photonic weightbank and integrated with an array of capacitive DEOAM on the 90 nm GF9WG monolithic process. The MNIST dataset serves as a benchmark for training and inference and reveals tradeoffs in neural network design [24]. Our approach to integrating analog memory in neuromorphic photonic processors aims to create a flexible processor with high bandwidth, efficient training capabilities, and the ability to adapt through on-chip, online learning.
1.2 Dynamic Electro-Optic Analog Memory (DEOAM)
@@
In a neuromorphic photonic processor with n rows and n columns of MRRs acting as weights, DEOAM reduces the number of DACs, thereby reducing power consumption. Since the conventional SRAM-DAC implementation requires each MRR to be paired with a DAC, scaling the implementation results in the number of DACs increasing by n2 (Figure 1b). The power consumed by SRAM-DAC is calculated as:
$P_{SRAM-DAC}=n^{2}P_{DAC}+P_{SRAM_{dynamic}}n+P_{SRAM_{static}}n^{2}$ (1)
where PSRAM-DAC is the total SRAM-DAC array power consumption in W, and PSRAM and PDAC are the SRAM and DAC power consumption in W, respectively, which includes static and dynamic power consumption and accounts for varying bit precisions of the DAC. All surveyed DACs are demonstrated, and the SRAMs associated with each DAC are surveyed to be in the same process node [12, 13, 14, 15, 16, 17, 18, 19, 20].
In contrast, using DEOAM reduces the number of DACs below n2 (Figure 1a) because data is held directly on the photonic device via analog memory, eliminating the constant processing typically needed by DACs. This concept is similar to the functionality in modern liquid crystal active-matrix displays [25]. The power consumed by DEOAM is calculated as:
$P_{DEOAM}=(P_{DAC}+\alpha CfV^{2})n$ (2)
where PDEOAM is the total DEOAM array power consumption, $\alpha$ is the activity factor (assumed to be 1), C is the load capacitance in F, f is the switching frequency in Hz and is half the sampling rate, and V is the switching voltage. DACs are still required in the DEOAM architecture to load initial weights from non-volatile, digital memory before beginning inference or training and to refresh analog memory during inference. However, this approach not only simplifies control complexities by reducing the number of electrical lines needed for DAC operations but also, with longer retention times, diminishes the need for DACs to continuously drive the photonic devices, leading to further power savings as the processor scales (Figure 1c, inset). Hence, integrating DEOAM and reducing the reliance on DACs paves the way for more efficient computing.
@@ 2.1 Electro-Optic Characteristics
2.2 Analysis of Analog Memory Specifications
To verify different analog memory specifications within a neural network system, a weight bank architecture, depicted in Figure 4, is emulated to assess analog memory performance during inference and training using the MNIST dataset [24].
Figure 4: The neural network emulation architecture and initial training. a) The feedforward neural network is trained using the MNIST dataset with 50000, 1-bit, 28x28 pixel images using a batch size of 64, then validated and tested using 10000 images [24]. The neural network is three layers with the input layer supporting 784 values for the 28x28 pixel images, the hidden layer supporting 50 neurons using a ReLU activation function, and the output layer supporting 10 neurons using a logarithmic softmax activation function to classify as a number. The neural network can map to the proposed neuromorphic photonic hardware that consists of 10 photonic cores with 50 rows of weight banks and 80 MRRs in each weight bank. Each core uses 80 wavelengths and 50 semiconductor optical amplifiers (SOA) to compensate for the splitting loss. b) The neural network achieves more than 95% inference accuracy after one epoch for all numbers. The specifics of the neural network architecture, performance, and modeling are detailed in Appendix F.
In Figure 5, the weight bank array is characterized for inference accuracy by comparing inference accuracy using weights trained on a GPU with full floating point accuracy and weights trained on emulated hardware with non-idealities in the analog memory and hardware. These non-idealities include analog memory control bit precision, noise, latency, and retention time, which can be generalized to any analog memory technology. These non-idealities are examined individually to assess their impact on inference accuracy. Figure 5a reveals that to achieve inference accuracies above 80%, more than 3 control bits are required for analog memory when weights are trimmed only in inference. The minimum control bit requirement increases to over 5 control bits when weights are adjusted in training and inference because lower control bit precision during training results in less accurate gradient calculations and weights, leading to reduced inference accuracy.
@@ 2.2 Analysis of Analog Memory Specifications
Once weights become unusable due to leakage, dead-time is consumed to refresh the weight, which can be approximated by the analog memory’s write time. When examining the ratio of retention time to write time for analog memories, this ratio (10,000 for DEOAM) is a couple orders of magnitude greater than the retention time to network latency ratio (100) required to have high inference accuracy. This means the dead-time for refreshing can be much shorter than network latency. Therefore, refreshes can occur after each image classification, thereby maintaining accurate weights for high inference accuracy without adding significant latency but at the cost of refresh power.
Figure 5: Neural network emulation results with and without non-idealities from the analog memory and system, such as control bit precision, noise, leakage, and latency. a) Minimum analog memory control bit precision to achieve more than 95% inference accuracy is 4 bits for weights trimmed only in inference and 8 bits for weights trimmed in inference and training. Inference accuracy degrades with increasing laser, SOA, PD, and TIA noise, but inference accuracy degrades b) earlier when weights are trained without noise and c) later when weights are trained with noise. d) The sweep of network latency and retention time characterizes inference accuracy for weights trained with and without leaky analog memory and network latency. Network latency and analog memory retention time constant at the point when inference accuracy degrades by 10% is plotted. When weights are trained with the effects of network latency and memory retention time, the network compensates for these effects and allows for a lower analog memory retention time constant requirement. Batch sizes of 64, 32, and 16 reveal that large batch sizes require longer retention times since weights are updated less frequently. Weights are updated more frequently in smaller batch sizes. Appendix F expands on results in d).
@@ 3 Discussion
However, scaling the neuromorphic photonic processor with analog memory presents challenges in analog memory and MRR control. Depending on the analog memory requirements, specialized circuitry may be necessary for writing to or refreshing the memory. As the system scales, strategies like sharing drive lines (as illustrated in Figure 1a) where DACs are shared along columns, can be essential in reducing power consumption. Furthermore, as the network grows with more weights, memory refresh rate becomes limited to DAC speeds and the size of the network. Accurate resonance control in MRRs is vital for maintaining precise optical weight representation. As the processor scales, implementing thermal control and multi-MRR control algorithms becomes increasingly critical for system stability [44, 45, 46, 47, 39, 48].
@@ 4 Conclusion
We demonstrated a proof-of-concept neuromorphic photonic circuit with DEOAM and verified its capabilities in a neuromorphic system. DEOAM reduces DACs in a neuromorphic photonic processor from n2 to n, thereby alleviating bottlenecks in bandwidth and energy consumption. By integrating analog memory on chip, neuromorphic photonic processors have the potential to achieve fast and efficient training. Since the training process occurs on chip, the neural network can readily adapt to environmental fluctuations and noise by incorporating them during training. Moreover, analog memory with on-chip backpropagation hardware can enable online training, allowing weights to be updated after each new input data. This opens up possibilities for various applications requiring recurrent events, event-based decision-making, or adaptability, such as time series forecasting. As a result, the next generation of AI processors, equipped with integrated analog memory, can help sustain the continued growth of AI.
@@ Appendix A Hardware Setup
The write time measurement setup is shown in Figure 7b and is similar to the retention time and leakage measurement setup in Figure 7a and 7c, respectively, except that the AWG is driving the analog memory and PN junction MRR while the SMU is holding the EN signal active. The write time measurement sequence is shown in Figure 3c and 3e. The AWG sends a pulse train with 5 ns edge rates to drive the analog memory and PN junction MRR, which modulates the optical response on the THRU and DROP port. The optical response is detected using the THRU port by the PD255 and measured by the DSO81304A oscilloscope. The write time measurement sequence shown in Figure 3d differs from Figure 3c and 3e because the rather than driving the analog memory and PN junction with the 5 ns edge rate pulse train, the transmission gate is driven with a pulse train that connects and disconnects the voltage source from the analog memory and PN junction. The rise time, fall time, and one time constant values are extracted from the oscilloscope time domain response.
@@ Appendix B Optical Characteristics
From the static optical characteristics of the MRRs in Figure 8a and 8b, the minimum wavelength spacing between different rings is 0.34 nm, the free spectral range (FSR) is 4.67 nm, the extinction ratio is between 15.6 dB and 19.6 dB, and the quality factor is about 50,000. From Figure 8a, optical bus power is estimated to be -6 dBm. In Figure 8c and Figure 8d, to observe the tuning efficiency of the PN junction MRRs, one MRR is biased at various voltages from 0 to 3 V (3 V is the limit for the thick oxide transistors) while the other MRRs are biased at zero voltage. The resonance wavelength and optical power of each MRR are observed at each voltage bias point to track tuning. Figure 8c reveals that as peak 3 is biased, peaks 1, 2, and 4 track similar trajectories, meaning optical crosstalk between adjacent resonances is negligible. In Figure 8d, the resonance power shift is less apparent than the resonance wavelength shift because the power shift occurs near the resonance of the ring, where the ring response’s slope is shallow. Looking at the tuning efficiency of the MRR in Figure 8e and 8f, the wavelength and power shifts are 6.22 pm/V and 0.23 dB/V, respectively.
@@ Appendix C Electro-Optic Characteristics
Appendix D Device-Level Comparison of Analog Memory Technologies
[…]
Appendix E System-Level Comparison of Analog Memory Technologies
Appendix F Emulation Modeling
F.1 Neural Network Architecture and Performance
@@ F.2 Noise
F.3 Weight Decay and Network Latency
@@
F.4 Effects of Retention Time Constant, Network Latency, and Batch Size on Inference Accuracy